# Controlling Large Language Model-based Agents for Large-Scale Decision-Making: An Actor-Critic Approach

### 09 Dec 2023 v2

@@ Abstract
The significant advancements in large language models (LLMs) have presented novel opportunities for tackling planning and decision-making within multi-agent systems. However, as the number of agents increases, the issues of hallucination in LLMs and coordination in multi-agent systems (MAS) have become increasingly pronounced. Additionally, the efficient utilization of tokens becomes a critical consideration when employing LLMs to facilitate the interactions of large numbers of agents. In this paper, we present a novel framework aimed at enhancing coordination and decision-making capabilities of LLMs within large-scale multi-agent environments. Our approach draws inspiration from the actor-critic framework employed in multi-agent reinforcement learning, and we develop a modular and token-efficient solution that effectively addresses challenges presented by LLMs and MAS. Through evaluations conducted in experiments involving system resource allocation and robot grid transportation, we demonstrate the considerable advantages afforded by our proposed approach.
Controlling Large Language Model-based Agents for Large-Scale Decision-Making: An Actor-Critic Approach
[…]
@@ 1 Introduction
Relying on training from massive datasets to capture extensive common knowledge and having demonstrated certain reasoning capabilities, Large Language Models (LLMs) have been widely applied and explored across various domains, rapidly emerging as powerful tools Brown et al. (2020); Kojima et al. (2022); Ruan et al. (2023a); Yang et al. (2023). Examples of such applications include question-answering systems Mallen et al. (2023), common-sense reasoning Hao et al. (2023), programming Tian et al. (2023), and embodied intelligence Driess et al. (2023). Recently, in the fields of natural language processing (NLP) and multi-agent systems (MAS), numerous research endeavors are dedicated to exploring the collaborative accomplishment of tasks through the utilization of multiple LLM-based agents. These efforts leverage the methodologies of role-playing Li et al. (2023) and debate Chan et al. (2023) as means to facilitate synergy and effective coordination among the agents involved. However, the majority of existing works primarily concentrate on coordination among a limited number of agents. The application of LLMs for effective coordination in large-scale agent scenarios has been rarely discussed.
The emergence of LLM-based AI agents has opened up new possibilities for addressing collaborative problems in MAS Zhang et al. (2023d, c). The introduction of prompting techniques, such as chain-of-thought (CoT) Wei et al. (2022), has significantly enhanced the reasoning and planning capabilities of LLMs. It avoids the need for training from scratch by providing an acceptable initial strategy based on common knowledge. Despite these advancements, applying LLMs directly to large-scale multi-agent decision-making tasks still presents numerous challenges. (1) As the number of agents increases, the joint action space grows exponentially, amplifying the difficulty of exploration and exploitation in complex MAS. (2) The limitations of LLMs themselves, such as the issue of hallucinations Zhang et al. (2023e), can affect the reliability of decision-making. (3) Effectively managing tokens or communication resources poses a significant challenge in large-scale scenarios involving LLM-based agents, so the design of token-efficient solutions is important and should not be overlooked.
Table 1: Comprehensive comparison of multi LLM-based agets methods. All approaches rely on either multi-agent debate or role-playing to accomplish decision-making tasks and solve NLP problems (task solver), or simulate collective behavior (community simulator).
Type Method Target Role Agents Num.
@@
To this end, we present LLaMAC, a novel framework for achieving a comprehensive decision-making process in collaborative tasks involving large-scale LLM-based agents, drawing inspiration from the classical actor-critic reinforcement learning (RL) approach Konda and Tsitsiklis (1999). Within LLaMAC, we design a centralized critic which takes on the role of a coordinator, making suggestions to each actor based on their decision memory. Subsequently, the actors interact with the environment, receive assigned tasks, and execute corresponding actions. (1) To tackle the exploration-exploitation trade-off inherent in the decision-making process, we introduce a TripletCritic structure inspired by the distributional code for value in the brain Dabney et al. (2020). This structure coordinates multiple critics with the same objective but different preferences. (2) We also establish a comprehensive feedback mechanism that incorporates both internal feedback within the TripletCritic and external feedback between the LLM-based actors (i.e., agents) and the TripletCritic. This mechanism aims to mitigate hallucination issues and bolster the robustness of LLM decision-making. (3) To construct a token-efficient planning framework, we incorporate mechanisms such as redundant memory information deletion and external feedback judgment. These mechanisms effectively reduce the length of context tokens and minimize the frequency of LLM accesses.
We first evaluate the performance of our method on a system resource allocation task to demonstrate its ability to strike a balance between exploration and exploitation, as well as its capability in large-scale multi-agent decision-making tasks. We further deploy our method in a more complex robot grid transportation scenario to validate its planning and decision-making capabilities. Experimental results demonstrate that our method outperforms existing approaches in terms of final performance, token utilization efficiency, and policy stability. To the best of our knowledge, we are the first to apply LLMs to large-scale multi-agent decision-making tasks.
@@ 2.2 Planning and Reasoning with LLM
Learning in massive corpora gives LLMs certain commonsense reasoning capabilities Kojima et al. (2022). Although there are still challenges in solving complex decision tasks, a large amount of work has proven that their methods can effectively improve the planning ability of LLMs Zelikman et al. (2022); Creswell et al. (2022). One line of research focuses on decomposing complex queries into sequential intermediate steps, known as Chain-of-Thought (CoT) Wei et al. (2022), to achieve accurate solutions. Another direction involves incorporating feedback mechanisms, showcasing their extensive capabilities in tackling complex decision-making challenges Wang et al. (2023b). Moreover, recent studies have begun to address this issue employing multiple LLMs. These approached are enhanced in their planning capabilities through techniques such as debate Chan et al. (2023); Liang et al. (2023) or role-playing Li et al. (2023); Hong et al. (2023). In the domain of decision-making, a subset of research utilizes prompting techniques to construct comprehensive processes covering perception, planning, and action, in cluding video games Zhang et al. (2023c), robot contro l Zhang et al. (2023d); Mandi et al. (2023), and open-world tasks Zhu et al. (2023); Gong et al. (2023). There are some studies about task planning and external tool usage Ruan et al. (2023b); Kong et al. (2023). However, as shown in Table 1, nearly all existing studies have only focused on tasks involving a small number of agents, with the involvement of a larger number of agents only observed in methods analyzing community simulations that do not require task-solving. However, it is worth noting that the existing studies, as outlined in Table 1, have predominantly concentrated on tasks involving a limited number of agents. The involvement of a larger number of agents has primarily been observed in methods that analyze community simulations, wherein task-solving is not a requirement. In light of this observation, our work uniquely emphasizes the application of language models in the realm of decision-making within large-scale multi-agent systems.
3 Method
@@ 3.1 Problem Formulation
This study focuses on the collaborative task planning of MAS, which can be formalized as a Goal-Augmented Decentralized Partially Observable Markov Decision Process (GA-Dec-POMDP) Spaan (2012). It is defined by a tuple: $\Gamma\triangleq\langle\mathcal{I},\mathcal{S},\mathcal{G},\{\mathcal{O}^{i}\}_{i\in\mathcal{I}},\{\mathcal{A}^{i}\}_{i\in\mathcal{I}},\mathcal{P},R\rangle$, where $\mathcal{I}$, $\mathcal{S}$, $\mathcal{G}$, $\mathcal{O}$, and $\mathcal{A}$ represent the sets of agents, state space, goal space, observation space, and action space, respectively. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the dynamic transition function, and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function. Within a given state $s\in\mathcal{S}$, each agent $i\in\mathcal{I}$ possesses its own local observations $o^{i}\in\mathcal{O}^{i}$ within its field of view and performs action $a^{i}\in\mathcal{A}^{i}$ accordingly. Formally, this problem requires each agent $i$ to learn a decision policy $\pi^{i}:\mathcal{O}^{i}\rightarrow\mathcal{A}^{i}$ to solve the task with a goal, which is equivalent to maximizing cumulative rewards.
3.2 LLaMAC: LLM-guided Collaborative Large-Scale Decision-Making
In this section, we present a systematic and modular formulation designed for LLM-based agents, namely Large Language Model-based Actor-Critic (LLaMAC), with a specific emphasis on their suitability for large-scale decision-making contexts. As illustrated in Figure 1, the overall framework adopts the Centralized Critic with Decentralized Actor (CCDA) structure, where actors and critics are LLM-based agents. The system incorporates three fundamental modules to facilitate a comprehensive decision-making process, enabling iterative reasoning, planning, and continuous interaction between the agents and the environment. These modules include an execution module responsible for action implementation and environment interaction, a memory module that stores information related to short-term experiences and long-term events, and a critic module that evaluates events and serves as a central coordinator. The functionalities of each module are as follows:
@@ 3.2 LLaMAC: LLM-guided Collaborative Large-Scale Decision-Making
Memory Module. The memory module serves to store crucial information needed during the decision-making process to aid the accumulation of useful knowledge and enhance the agent’s decision-making capabilities. Specifically, the short-term memory is used to store the most recent state. In contrast, the historical trajectory and experiential information learned from interactions are stored in the long-term memory. The memory module also incorporates a mechanism for filtering redundant information. During long-term planning processes, it retains only the most recent L steps of state transitions $<s_{t-L+1},a_{t-L+1},r_{t-L+1},s_{t-L+2},...,s_{t}>$. This assists the agent in comprehending the relationship between actions and changes in environmental states. Additionally, in scenarios involving single-step decision-making, repeated state-action pairs are not stored.
Critic Module. The critic module assumes a central role within the workflow of LLaMAC, exerting significant influence. It receives the present state and extracts pertinent details from the memory module, enabling evaluation and learning from the actors’ historical trajectories. Functioning as a centralized coordinator, the critic module engages in reasoning and planning activities to formulate potential high-reward and reliable plan suggestions. These suggestions then serve as guides for the interaction between the actor and the environment.
@@
    The critic receives the state and extracts the relevant information from the memory module. Utilizing these inputs, it facilitates a three-critic dialogue and subsequently generates the textual suggestion denoted as $su$ for each actor. The relevant memory is initially empty during the first interaction, and the specific methodology for generating the suggestion is elaborated upon in Section 3.2.1.
@@
    Each actor is provided with the observation denoted as $o$ from the environment, as well as the suggestion $su$ from the TripletCritic. Subsequently, actors engage in a process called Plan Confirmation (refer to Section 3.2.2 for comprehensive information on this process).
@@
3.2.1 TripletCritic with Internal Feedback
Figure 2: Internal Feedback within the TripletCritic.
@@ 3.2.1 TripletCritic with Internal Feedback
To this end, we develop the TripletCritic, which incorporates an internal feedback mechanism. The design of TripletCritic is inspired by the distributed encoding of reward and value by dopamine neurons in the brain Dabney et al. (2020). Each dopamine neuron contains partial information about the actual reward, and different dopamine neurons utilize different value predictions, enabling the brain to model the value distribution of reward events. Similarly, as depicted in Figure 2, the TripletCritic framework encompasses a dual-critic structure, each with the same objective but distinct preferences, alongside the third critic, called the assessor, who assumes the responsibility of reconciling these preferences. One critic exhibits a proclivity for exploration, prioritizing long-term gains, while the other gravitates towards exploitation, emphasizing short-term gains. The assessor fulfills two primary roles. Firstly, it makes Veracity Scrutiny to check the strategies employed by the dual-critic, offering internal feedback in the event of errors. Secondly, it undertakes Belief Correction in order to establish a harmonious equilibrium between exploration and exploitation within the planners. Additionally, the assessor collaborates with the actor to transmit the final suggestion assignment, informed by these assessments and corrections.
By employing an internal feedback mechanism and an evaluation mechanism that trades off different preferences, the TripletCritic effectively mitigates the occurrence of hallucination issue. This approach enhances the feasibility and robustness of the initial policy and further facilitates the development of token-efficient solutions.
3.2.2 External Feedback from Actor to Critic
Figure 3: External feedback mechanism from actor to critic.
@@ 3.2.2 External Feedback from Actor to Critic
The TripletCritic provides each actor with a potential initial feasible solution. To facilitate the iterative long-term planning process and achieve the ultimate goal, as well as to reduce the access costs of decision-making for a large number of intelligent agents, we have additionally incorporated an external feedback mechanism from actor to critic.
Initially, as depicted in Figure 3, the TripletCritic sends suggestions $\{su^{i}\}_{i\in\mathcal{I}}$ to each actor, and all actors pass the proposed plans through an external Plan Confirmation to determine their feasibility. If further improvements are deemed necessary, the corresponding LLM is accessed. The LLM takes as input the agent’s observation $o^{i}$ and the corresponding suggestion $su^{i}$, providing insights into the underlying issues and potential enhancement strategies. Once feedback is received from all actors, the information is aggregated and sent back to the Assessor within the TripletCritic. The Assessor utilizes the internal feedback dialogue information and the actors’ external feedback to further update the suggestions for actors with identified issues, returning new suggestions to the respective actors. This iterative process continues until all actors determine that no further improvements are necessary, at which point actions are executed directly. It is worth noting that due to the reliability of the TripletCritic, the opportunity for actors to provide external feedback is limited, resulting in minimal access costs. The occasional feedback process further enhances the performance of the final strategy.

### 23 Jan 2024 v3

@@ Abstract
The remarkable progress in Large Language Models (LLMs) opens up new avenues for addressing planning and decision-making problems in Multi-Agent Systems (MAS). However, as the number of agents increases, the issues of hallucination in LLMs and coordination in MAS have become increasingly prominent. Additionally, the efficient utilization of tokens emerges as a critical consideration when employing LLMs to facilitate the interactions among a substantial number of agents. In this paper, we develop a modular framework called LLaMAC to mitigate these challenges. LLaMAC implements a value distribution encoding similar to that found in the human brain, utilizing internal and external feedback mechanisms to facilitate collaboration and iterative reasoning among its modules. Through evaluations involving system resource allocation and robot grid transportation, we demonstrate the considerable advantages afforded by our proposed approach.
@@ 1 Introduction
Relying on training from massive datasets to capture extensive common knowledge and having demonstrated certain reasoning capabilities, Large Language Models (LLMs) have been widely applied and explored across various domains, rapidly emerging as powerful tools Brown et al. (2020); Kojima et al. (2022); Ruan et al. (2023a); Yang et al. (2023). The utilization of prompting techniques, such as chain-of-thought (CoT) Wei et al. (2022), has played a pivotal role in further augmenting the reasoning and planning capabilities of LLMs. This approach eliminates the need for training from scratch by providing an acceptable initial strategy based on common knowledge. Examples of such applications include question-answering systems Mallen et al. (2023), common-sense reasoning Hao et al. (2023), programming Tian et al. (2023), and embodied intelligence Driess et al. (2023).
Recently, in the fields of natural language processing (NLP) and multi-agent systems (MAS), numerous research endeavors are dedicated to exploring the collaborative task-solving potential facilitated by the cooperation of multiple agents grounded in LLMs. These efforts leverage the role-playing Li et al. (2023) and debate Chan et al. (2023) to facilitate synergy and effective coordination among the agents involved. However, most existing works focus on coordinating a limited number of agents as shown in Table 1. The application of LLMs for effective coordination in large-scale agent scenarios has received limited attention. This is attributed mainly to the significant increase in complexity and difficulty associated with applying LLMs to large-scale multi-agent decision-making tasks.
In this paper, we direct our attention to the following key challenges: (1) As the number of agents increases, the joint action space grows exponentially, amplifying the difficulty of exploration and exploitation in complex MAS. (2) The limitations of LLMs themselves, as highlighted by the issue of hallucinations Zhang et al. (2023e), can affect the reliability of decision-making. (3) Effectively managing tokens or communication resources presents a substantial challenge in scenarios involving large-scale LLM-based agents. We prioritize these challenges due to their inherent and widespread nature in large-scale settings, and the absence of comprehensive solutions. By focusing on these general challenges, we try to contribute insights and solutions that hold broad relevance, offering a foundational framework for tackling intricacies in diverse real-world scenarios.
Table 1: Comprehensive comparison of LLM-based multi-agent methods. All approaches rely on either multi-agent debate or role-playing to accomplish decision-making tasks and solve NLP problems (task solver), or simulate collective behavior (community simulator).
Type Method Target Agent Configuration Agents Num.
@@
To this end, we present Large Language Model-based Actor-Critic (LLaMAC), a novel framework for achieving a comprehensive decision-making process in collaborative tasks involving large-scale LLM-based agents, drawing inspiration from the classical actor-critic reinforcement learning (RL) approach Konda and Tsitsiklis (1999). Within LLaMAC, we design a centralized critic which takes on the role of a coordinator, making suggestions to each actor based on their decision memory. Subsequently, the actors engage in interactions with the environment, receiving assigned tasks, conducting analyses, and performing corresponding actions. Specifically, our primary contributions are as follows:
  • To attain a viable and robust initial strategy and tackle the exploration-exploitation trade-off inherent in the decision-making process, we introduce the TripletCritic structure, which is inspired by the distributional code for value in the brain Dabney et al. (2020). This architecture effectively coordinates multiple critics with shared objectives but varying preferences through internal feedback, thereby providing dependable action suggestion for each actor involved.
  • We also establish an external feedback mechanism between the LLM-based actors (i.e., agents) and the TripletCritic. This mechanism serves to not only reduce the access cost of the LLM but also allows each actor to maintain independent exploration and decision-making capabilities.
  • We propose a modular and token-efficient framework for augmenting the decision-making capabilities of LLM-based agents in large-scale multi-agent environments. This framework enables autonomous iterative cooperation among a large number of agents.
We first evaluate the performance of our method on a system resource allocation task to demonstrate its ability to strike a balance between exploration and exploitation, as well as its capability in large-scale multi-agent decision-making tasks. We further deploy our method in a more complex robot grid transportation scenario to validate its planning and decision-making capabilities. Experimental results demonstrate that our method outperforms existing approaches in terms of final performance, token utilization efficiency, and policy stability. To the best of our knowledge, we are the first to apply LLMs to large-scale multi-agent decision-making tasks involving more than 50 agents, as indicated in Table 1.
@@ 2.2 Planning and Reasoning with LLM
Learning in massive corpora gives LLMs certain commonsense reasoning capabilities Kojima et al. (2022). Although there are still challenges in solving complex decision tasks, a large amount of work has proven that their methods can effectively improve the planning ability of LLMs Zelikman et al. (2022); Creswell et al. (2022). One line of research focuses on decomposing complex queries into sequential intermediate steps, known as Chain-of-Thought (CoT) Wei et al. (2022), to achieve accurate solutions. Another direction involves incorporating feedback mechanisms, showcasing their extensive capabilities in tackling complex decision-making challenges Wang et al. (2023b). Moreover, recent studies have begun to address this issue employing multiple LLMs. These approached are enhanced in their planning capabilities through techniques such as debate Chan et al. (2023); Liang et al. (2023) or role-playing Li et al. (2023); Hong et al. (2023). In the domain of decision-making, a subset of research utilizes prompting techniques to construct comprehensive processes covering perception, planning, and action, in cluding video games Zhang et al. (2023c), robot contro l Zhang et al. (2023d); Mandi et al. (2023), and open-world tasks Zhu et al. (2023); Gong et al. (2023). There are some studies about task planning and external tool usage Ruan et al. (2023b); Kong et al. (2023). However, it is worth noting that the existing studies, as outlined in Table 1, have predominantly concentrated on tasks involving a limited number of agents. The involvement of a larger number of agents has primarily been observed in methods that analyze community simulations, wherein task-solving is not a requirement. In light of this observation, our work uniquely emphasizes the application of language models in the realm of decision-making within large-scale multi-agent systems.
3 LLaMAC
In this section, we formally present a systematic and modular framework designed for LLM-based agents, namely Large Language Model-based Actor-Critic (LLaMAC), with a specific emphasis on their suitability for large-scale decision-making contexts.
@@ 3.1 Problem Formulation
This study focuses on the collaborative task solving of MAS, which can be formalized as a Goal-Augmented Decentralized Partially Observable Markov Decision Process (GA-Dec-POMDP) Spaan (2012). It is defined by a tuple: $\Gamma\triangleq\langle\mathcal{I},\mathcal{S},\mathcal{G},\{\mathcal{O}^{i}\}_{i\in\mathcal{I}},\{\mathcal{A}^{i}\}_{i\in\mathcal{I}},\mathcal{P},R\rangle$, where $\mathcal{I}$, $\mathcal{S}$, $\mathcal{G}$, $\mathcal{O}$, and $\mathcal{A}$ represent the sets of agents, state space, goal space, observation space, and action space, respectively. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the dynamic transition function, and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents the reward function. Within a given state $s\in\mathcal{S}$, each agent $i\in\mathcal{I}$ possesses its own local observations $o^{i}\in\mathcal{O}^{i}$ within its field of view and performs action $a^{i}\in\mathcal{A}^{i}$ accordingly. Formally, this problem requires each agent $i$ to learn a decision policy $\pi^{i}:\mathcal{O}^{i}\rightarrow\mathcal{A}^{i}$ to solve the task with a goal, which is equivalent to maximizing cumulative rewards.
3.2 Overall Framework
As illustrated in Figure 1, LLaMAC introduces the Centralized Critic with Decentralized Actor (CCDA) structure, where actors and critics are LLM-based agents. The system incorporates three fundamental modules to facilitate a comprehensive decision-making process, enabling iterative reasoning, planning, and continuous interaction between the agents and the environment. The functionalities of each module are as follows:
Figure 2: Internal Feedback within the TripletCritic (Left) and External Feedback mechanism from actor to critic (Right).
@@ 3.2 Overall Framework
Memory Module. The memory module serves to store crucial information needed during the decision-making process to aid the accumulation of useful knowledge and enhance the agent’s decision-making capabilities. Specifically, the short-term memory is used to store the most recent state. In contrast, the historical trajectory and experiential information learned from interactions are stored in the long-term memory. The memory module also incorporates a mechanism for filtering redundant information. During long-term planning processes, it retains only the most recent $L$ steps of state transitions $<s_{t-L+1},a_{t-L+1},r_{t-L+1},s_{t-L+2},...,s_{t}>$. This assists the agent in comprehending the relationship between actions and changes in environmental states.
Critic Module. The critic module assumes a central role within the workflow of LLaMAC. It receives the present state and extracts pertinent details from the memory module, enabling evaluation and learning from the actors’ historical trajectories. Functioning as a centralized coordinator, the critic module engages in reasoning and planning activities to formulate potential high-reward and reliable plan suggestions. These suggestions then serve as guides for the interaction between the actor and the environment.
@@
    The critic receives the state and extracts the relevant information from the memory module. Utilizing these inputs, it facilitates a three-critic dialogue (Internal Feedback) and subsequently generates the textual suggestion denoted as $su$ for each actor.
@@
    Each actor is provided with the observation denoted as $o$ from the environment, as well as the suggestion $su$ from the TripletCritic. Subsequently, actors engage in a process called External Feedback.
@@
3.3 TripletCritic with Internal Feedback
@@
To this end, we develop the TripletCritic, which incorporates an internal feedback mechanism. The design of TripletCritic is inspired by the distributed encoding of reward and value by dopamine neurons in the brain Dabney et al. (2020). Each dopamine neuron contains partial information about the actual reward, and different dopamine neurons utilize different value predictions, enabling the brain to model the value distribution of reward events. Similarly, as depicted in Figure 2, the TripletCritic framework encompasses a dual-critic structure, each with the same objective but distinct preferences, alongside the third critic, called the assessor, who assumes the responsibility of reconciling these preferences. One critic exhibits a proclivity for exploration, prioritizing long-term gains, while the other gravitates towards exploitation, emphasizing short-term gains. The assessor fulfills two primary roles. Firstly, it makes Veracity Scrutiny to check the strategies employed by the dual-critic, offering internal feedback in the event of errors. Secondly, it undertakes Belief Correction in order to establish a harmonious equilibrium between exploration and exploitation within the planners. Additionally, the assessor collaborates with the actors to transmit the final suggestion assignment, informed by these assessments and corrections.
3.4 External Feedback from Actor to Critic
@@
The TripletCritic provides each actor with a potential initial feasible solution. To facilitate the iterative long-term planning process and achieve the ultimate goal, as well as to reduce the access costs of decision-making for a large number of intelligent agents, we additionally incorporate an external feedback mechanism from actor to critic.
Initially, as depicted in Figure 2, the TripletCritic sends suggestions $\{su^{i}\}_{i\in\mathcal{I}}$ to each actor, and all actors pass the proposed plans through an external Plan Confirmation to determine their feasibility. If further improvements are deemed necessary, the corresponding LLM is accessed. The LLM takes as input the agent’s observation $o^{i}$ and the corresponding suggestion $su^{i}$, providing insights into the underlying issues and potential enhancement strategies. Once feedback is received from all actors, the information is aggregated and sent back to the Assessor within the TripletCritic. The Assessor utilizes the internal feedback dialogue information and the actors’ external feedback to further update the suggestions for actors with identified issues, returning new suggestions to the respective actors. This iterative process continues until all actors determine that no further improvements are necessary, at which point actions are executed directly.
The coordination among various modules is facilitated by both internal and external feedback, thus forming a comprehensive and automated iterative planning process. TripletCritic enhances the viability and robustness of the initial policy by incorporating an internal feedback mechanism and an evaluation mechanism that balances different preferences. Additionally, it effectively reduces the occurrence of hallucination issues. It is important to highlight that the reliability of TripletCritic reduces the actors’ opportunity to provide external feedback, thereby minimizing access costs and promoting the development of token-efficient solutions. The occasional external feedback process further improves the performance of the ultimate strategy.